Zinc reacts with hydrochloric acid according to the reaction equation Zn(s) +2HCl(aq) ? ZnCl2(aq)-H2(g How many milliliters of 4.50 M HCl(aq) are required to react with 5.25 g of an ore containing 50.0% Zn(s) by mass? Number mLplease help
Answer #1

Amount of Zn in the ore = 0.50*5.25g = 2.625 g

Molecular weight of Zn = 65 g/mol

Moles of Zn in the ore = 2.625/65 = 0.0403

One mole of Zn consumes 2 mole of HCl

Therefore, 0.0403 moles of Zn will consume 0.0806 moles of HCl

Now, Molarity = moles/Volume (in L)

4.50 = 0.0806/Volume

Volume = 0.0179 L = 17.9 mL